# Indoor Scene Reconstruction with Fine-Grained Details Using Hybrid Representation and Normal Prior Enhancement

### v2

@@ 1 Introduction
Encoding scenes with a neural implicit function [1, 2, 3] has attracted increasing attention due to its compactness and good performance. Some recent works have achieved remarkable results in 3D scene reconstruction from 2D supervision, by combining neural implicit function and differentiable volume rendering. NeuS [4] and VolSDF [5] can reconstruct high-quality watertight surfaces of a single object by representing the geometry as an implicit signed distance field (SDF). Follow-up works try to integrate the implicit SDF representation and additional geometric priors, including semantic segmentation [6], monocular depth [7], or surface normals [8], to address the challenge of indoor scene reconstruction. Compared to traditional MVS-based [9, 10] or TSDF-based [11, 12] approaches, these neural implicit methods can produce complete and promising surfaces, but still struggle with delicate and complex structures.
@@
As for the enhancement of predicted normal priors, we observe that artifacts present in the input RGB images can greatly affect the quality of predicted normals. Thus, we devise a sharpening and denoising technique before feeding the images into the prior estimation module, which effectively reduces the errors in predicted normals. Furthermore, we design an uncertainty module, which predicts the pixel-wise uncertainty maps of the estimated normal priors. Besides the RGB images and monocular estimated normals, this module also takes features from the visual foundation model [20] as input to supplement the prediction with high-level semantic information. We then treat the predicted uncertainty as the weight of normal prior supervision. This uncertainty module determines the reliability of normal priors and reduces the negative impact of incorrectly predicted priors on reconstruction quality. In conclusion, we summarize our contributions as follows:
@@ 2.1 Neural Surface Representation
Representing geometric surfaces by neural implicit functions has recently received increasing attention, because of its compactness and remarkable performance. The seminal work of DeepSDF [1] first proposes to use a neural network to model a signed distance field, which encodes the underlying geometry of the target object. However, DeepSDF requires ground-truth 3D meshes to supervise the learning process. Recent methods incorporate neural implicit functions and differentiable rendering to reconstruct surfaces only from the supervision of multi-view 2D images. DVR [21] first proposes a differentiable rendering formulation for implicit shape and texture representations. IDR [22] further designs an architecture that simultaneously learns the implicit SDF field, camera poses, and a neural renderer. UNISURF [23] utilizes occupancy to represent implicit surfaces and modifies the volume rendering equation correspondingly. Subsequently, NeuS [4] and VolSDF [5] define the density used in volume rendering process as logistic sigmoid function and Laplace’s cumulative distribution function applied to an SDF representation. Nevertheless, all those works use a single MLP to approximate the implicit function, which struggles to scale to large and complicated scenes due to the network capacity. There also exist other researches [24, 16, 15, 25, 26] that use multi-scale voxel grids followed by a shallow decoder to represent the scenes, but this leads to non-smoothness. Neuralangelo [27] leverages 3D hash grids to recover dense 3D surface structures. However, hash grids also suffer from localities and hash collisions. In this work, we devise a novel hybrid representation incorporating MLP and tri-plane to enhance the network’s ability to express high-frequency and delicate surfaces, while maintaining non-local smoothness.
@@ 2.2 Indoor Scene Reconstruction
Conventional multi-view stereo (MVS) methods [10, 9] often struggle to reconstruct large texture-less indoor scene areas densely. With the development of large-scale indoor datasets [28, 29], learning-based MVS methods are proposed to alleviate this problem. These methods [30, 31, 32, 33] usually predict a depth map for each frame and fuse these depth maps to build the final scene. However, the resulting surfaces tend to be noisy and incomplete due to the depth inconsistency problem. Another research branch [11, 34, 12, 35, 36] proposes directly regressing input images to truncated signed distance function (TSDF) volume. Then, the meshes can be extracted from TSDF using the marching-cube algorithm. Due to the limitation of TSDF volume resolution, the generated surfaces often lack details.
@@ 2.3 Uncertainty Estimation
Uncertainty estimation, which aims to quantify the reliability of predictions, can help recognize failure scenarios and enable robust applications. Uncertainty is crucial for many computer vision tasks such as optical flow [39], SLAM [40], and multi-view stereo [41]. Predicted uncertainty is used to facilitate the optimization through robust loss [42], guided sampling [43], outlier rejection [44], etc. However, the integration of the neural radiance field with uncertainty remains a relatively unexplored terrain. Some current efforts [45, 46, 47] mainly focus on the task of rendering, while we focus on surface reconstruction. In this work, we novelly leverage the estimated uncertainty in a weighted loss to guide our model to utilize more precise priors while avoiding misguidance by less accurate ones.
@@ 3.1 Preliminary
where $\Phi_{\tau}$ is the Sigmoid function with a learnable parameter $\tau$. Then, the difference between the rendered pixel color $C(\bm{r})$ and the ground-truth pixel color $\hat{C}(\bm{r})$ is minimized to optimize the $f_{g}$ and $f_{c}$. After optimization, the surface can be extracted as the zero level-set of the signed distance field $f_{g}$.
@@ 3.2 Hybrid Representation for Geometry
Empirically, we find that different areas of indoor scenes have different characteristics. The outer contours of a scene are usually smooth and flat, while the interior often contains delicate structures. Thus, in order to accurately reconstruct both the flat areas and the fine-grained details, we propose a hybrid geometry architecture, which contains an MLP branch and a tri-plane branch (see Figure 2). We aim to exploit the smoothness bias of MLP to encode the rough outline of the scene and the high-frequency bias of tri-plane features to encode the delicate structures. To encourage the two branches to focus on different areas of the scene, we find that the initialization plays a vital role. We initialize the MLP branch to produce an approximate SDF of a bounding sphere and the tri-plane branch to produce a zero SDF. In this way, we implicitly decouple an indoor scene into two distinct regions without the need for extra segmentation annotations. Given a 3D position $\bm{x}\in\mathbb{R}^{3}$, our pipeline first passes it through the MLP branch to get a coarse SDF value $\tilde{\bm{s}}$ and a coarse hidden feature $\tilde{\bm{h}}$. Then, the tri-plane branch refines these coarse outputs. Our tri-plane representation $T$ consists of three axis-aligned orthogonal feature planes. Each feature plane has the dimension of $N\times N\times C$, where $N$ is the spatial resolution, and $C$ is the number of channels. We retrieve the final feature $[T_{xy},T_{xz},T_{yz}]$ by projecting $\bm{x}$ onto these planes to find the tri-plane features and concatenating them together. A shallow, two-layer decoder further regresses the aggregated feature $[T_{xy},T_{xz},T_{yz}]$ to the residual SDF $\Delta\bm{s}$ and residual hidden feature $\Delta\bm{h}$. The final outputs are the summation of these two branches
@@ 3.3.2 Uncertainty Estimation Module
where $\mathcal{R}$ is the set of rays in each batch. $\bm{n}(\bm{r})$ and $\hat{\bm{n}}(\bm{r})$ are the rendered normal and the pseudo ground-truth normal, respectively. Specifically, for the uncertainty value corresponding to a ray $\bm{r}$, we normalize it to $[0,1]$, denoted as $\bm{u}(\bm{r})$. When $\mathcal{L}_{prior}$ supervises the implicit SDF field together with the typical color loss, our proposed uncertainty values can be viewed as a trade-off between the color information and normal information. For complicated regions with high uncertainty, the proposed loss reduces the weight of normal supervision, thereby strengthening the role of color supervision. By balancing the supervision weights, our model can resist being misled by unreliable normal priors and effectively utilize reliable normal priors, facilitating accurate and high-quality reconstruction.
To predict the pixel-wise uncertainty of the estimated normal maps, we design a novel module as shown in Figure 4. This module adopts a U-Net [49] architecture with RGB images and the estimated normal maps as input. To facilitate the uncertainty prediction, we also input the high-level DINO [20] features into the module, since the normal uncertainty is usually semantic-relevant. Semantic features are extracted from the last attention layer of the visual foundation model DINO, and then dimensionally reduced by PCA to retain the most important information. We train this uncertainty estimation module on a subset (disjoint from the test set) of the ScanNet dataset [28]. Concretely, we compute the difference between the estimated normals and normals rendered from real 3D meshes to obtain the ground-truth uncertainty maps, and use them to supervise our proposed estimation module.
@@ 3.4 Loss Functions
We optimize our model with the weighted sum of the normal prior loss $\mathcal{L}_{prior}$, the Eikonal loss $\mathcal{L}_{eik}$, and the RGB color loss $\mathcal{L}_{rgb}$:
@@ 4.3.1 Analysis of Geometry Representation
Figure 13: Results of the novel view synthesis task. In each block, the first row is the rendering results of our proposed method, and the second row is the corresponding ground-truth images.
@@ 4.3.2 Analysis of Appearance Embedding
Since we only supervise our model on 2D image observations, the inconsistency caused by photometric and environmental variations can lead to noisy geometries. Our experiments illustrate that the appearance embeddings can alleviate this problem, as the embeddings can learn to compensate for variations and guarantee a consistent reconstruction. We render the RGB images, normal maps, and depth maps in Figure 10. The depth maps and normal maps are sharper and cleaner by adding appearance embeddings.
@@ 4.3.3 Analysis of Prior Enhancement
As illustrated in Figure 10, imperfections like motion blur during the video capture can cause wrong predictions of normal priors (e.g., bottom of the piano), which hinder the reconstruction of fine and accurate geometries. After applying our proposed enhancement, the image becomes sharper, and the predicted normal maps get better. As a result, Table V shows that the accuracy and completeness of reconstructed meshes are improved as well.
@@ 4.3.4 Analysis of Uncertainty Estimation
The estimated uncertainty can serve as a measure of reliability and a balance between color supervision and normal supervision. In intricate areas, it guides our network to learn better geometry from rich color information rather than inaccurate normal information, thereby facilitating the correct reconstruction of exquisite structures. Ablation results show that the uncertainty estimation module helps to improve geometry quality. As fine-grained details take up only a small proportion of the scene and cannot be fully reflected by numerical metrics, we also visualize the reconstructed meshes and rendered normal maps in Figure 11. Adding uncertainty as guidance makes the reconstruction results more visually appealing and richer in high-frequency details. In addition, we find that supplementing the uncertainty estimation module with high-level semantic DINO features can further facilitate the reconstruction of complex and delicate structures (shown in Figure 14).
@@ 4.5 Novel View Synthesis
The novel view is another form of understanding 3D scenes apart from meshes. Although our primary goal is to reconstruct high-fidelity, detailed surfaces of an indoor scenario, we can also synthesize realistic images under novel views facilitated by the accurate reconstruction results. As we select one image as input every ten frames to train our model, we use the remaining views to conduct rendering. Figure 13 shows the rendering results of our approach and the corresponding ground-truth images.
@@ 4.6 Efficiency and Robustness
In Table VI, we compare the parameters and memory consumption of several baseline methods. ManhattanSDF [6] and MonoSDF [7] adopt the parameter compact MLP models but struggle to reconstruct fine geometries. The memory consumption and parameters of NueralAngelo [27] are large because it utilizes 3D hash grids, which can be regarded as a special version of voxel grids. HelixSurf [38] represents the implicit SDF field with a relatively small MLP accompanied by a PatchMatch-based MVS module. However, the joint optimization of MVS module and neural radiance field in HelixSurf consumes a large memory overhead. In contrast, our method can produce the best quality meshes with comparable parameters and less memory consumption.

### v3

@@ 1 Introduction
Encoding scenes with a neural implicit function [1, 2, 3] has attracted increasing attention due to its compactness and good performance. Some recent works have achieved remarkable results in 3D scene reconstruction from 2D supervision, by combining the neural implicit function and differentiable volume rendering. NeuS [4] and VolSDF [5] can reconstruct high-quality watertight surfaces of a single object by representing the geometry as an implicit signed distance field (SDF). Follow-up works try to integrate the implicit SDF representation and additional geometric priors, including semantic segmentation [6], monocular depth [7], or surface normals [8], to address the challenge of indoor scene reconstruction. Compared to traditional MVS-based [9, 10] or TSDF-based [11, 12] approaches, these neural implicit methods can produce complete and promising surfaces, but still struggle with delicate and complex structures.
@@
As for the enhancement of predicted normal priors, we observe that artifacts present in the input RGB images can greatly affect the quality of predicted normals. Thus, we devise a sharpening and denoising technique before feeding the images into the prior estimation module, which effectively reduces the errors in predicted normals. Furthermore, we design an uncertainty module, which predicts the pixel-wise uncertainty maps of the estimated normal priors. Besides the RGB images and monocular estimated normals, this module also takes features from the visual foundation model [20] as input to supplement the prediction with high-level structural information. We then treat the predicted uncertainty as the weight of normal prior supervision. This uncertainty module determines the reliability of normal priors and reduces the negative impact of incorrectly predicted priors on reconstruction quality. In conclusion, we summarize our contributions as follows:
@@ 2.1 Neural Surface Representation
Representing geometric surfaces by neural implicit functions has recently received increasing attention, because of its compactness and remarkable performance. The seminal work of DeepSDF [1] first proposes to use a neural network to model a signed distance field, which encodes the underlying geometry of the target object. However, DeepSDF requires ground-truth 3D meshes to supervise the learning process. Recent methods incorporate neural implicit functions and differentiable rendering to reconstruct surfaces only from the supervision of multi-view 2D images. DVR [21] first proposes a differentiable rendering formulation for implicit shape and texture representations. IDR [22] further designs an architecture that simultaneously learns the implicit SDF field, camera poses, and a neural renderer. UNISURF [23] utilizes occupancy to represent implicit surfaces and modifies the volume rendering equation correspondingly. Subsequently, NeuS [4] and VolSDF [5] define the density used in volume rendering process as logistic sigmoid function and Laplace’s cumulative distribution function applied to an SDF representation. Nevertheless, all those works use a single MLP to approximate the implicit function, which struggles to scale to large and complicated scenes due to the network capacity. There also exist other researches [16, 15, 24, 25] that use multi-scale voxel grids followed by a shallow decoder to represent the scenes, but this leads to non-smoothness. Neuralangelo [26] leverages 3D hash grids to recover dense 3D surface structures. However, hash grids also suffer from localities and hash collisions. In this work, we devise a novel hybrid representation incorporating MLP and tri-plane to enhance the network’s ability to express high-frequency and delicate surfaces, while maintaining non-local smoothness.
@@ 2.2 Indoor Scene Reconstruction
Conventional multi-view stereo (MVS) methods [10, 9] often struggle to reconstruct large texture-less indoor scene areas densely. With the development of large-scale indoor datasets [27, 28], learning-based MVS methods are proposed to alleviate this problem. These methods [29, 30, 31] usually predict a depth map for each frame and fuse these depth maps to build the final scene. However, the resulting surfaces tend to be noisy and incomplete due to the depth inconsistency problem. Another research branch [11, 32, 12, 33] proposes directly regressing input images to truncated signed distance function (TSDF) volume. Then, the meshes can be extracted from TSDF using the marching-cube algorithm. Due to the limitation of TSDF volume resolution, the generated surfaces often lack details.
@@ 2.3 Uncertainty Estimation
Uncertainty estimation, which aims to quantify the reliability of predictions, can help recognize failure scenarios and enable robust applications. Uncertainty is crucial for many computer vision tasks such as optical flow [36], SLAM [37], and multi-view stereo [38]. Predicted uncertainty is used to facilitate the optimization through robust loss [39], guided sampling [40], outlier rejection [41], etc. However, the integration of the neural radiance field with uncertainty remains a relatively unexplored terrain. Some current efforts [42, 43] mainly focus on the task of rendering, while we focus on surface reconstruction. In this work, we novelly leverage the estimated uncertainty in a weighted loss to guide our model to utilize more precise priors while avoiding misguidance by less accurate ones.
@@ 3.1 Preliminary
where $\Phi_{\tau}$ is the Sigmoid function with a learnable parameter $\tau$. Then, the difference between the rendered pixel color $C(\bm{r})$ and the ground-truth pixel color $\hat{C}(\bm{r})$ is minimized to optimize the $f_{g}$ and $f_{c}$. After optimization, the surface $\mathcal{S}$ can be extracted as the zero level-set of the signed distance field $\mathcal{S}=\{\bm{x}|f_{g}(\bm{x})=0\}$.
@@ 3.2 Hybrid Representation for Geometry
Empirically, we find that different areas of indoor scenes have different characteristics. The outer contours of a scene are usually smooth and flat, while the interior often contains delicate structures. Thus, in order to accurately reconstruct both the flat areas and the fine-grained details, we propose a hybrid geometry architecture, which contains an MLP branch and a tri-plane branch (see Figure 2). We aim to exploit the smoothness bias of MLP to encode the rough outline of the scene and the high-frequency bias of tri-plane features to encode the delicate structures. To encourage the two branches to focus on different areas of the scene, we find that the initialization plays a vital role. Following SAL [44], we initialize the MLP branch to produce approximate SDF of a unit bounding sphere, and initialize the tri-plane branch to produce a zero SDF. In this way, we implicitly decouple an indoor scene into two distinct regions without the need for extra segmentation annotations. Given a 3D position $\bm{x}\in\mathbb{R}^{3}$, our pipeline passes it through the MLP branch to get a coarse SDF $\tilde{\bm{s}}$ and a coarse hidden feature $\tilde{\bm{h}}$. Simultaneously, the tri-plane branch outputs the high-frequency complements. Our tri-plane representation $T$ consists of three axis-aligned orthogonal feature planes. Each feature plane has the dimension of $N\times N\times C$, where $N$ is the spatial resolution, and $C$ is the number of channels. We retrieve the final feature $[T_{xy},T_{xz},T_{yz}]$ by projecting $\bm{x}$ onto these planes to find the tri-plane features and concatenating them together. A shallow, two-layer decoder further regresses the aggregated feature $[T_{xy},T_{xz},T_{yz}]$ to the residual SDF $\Delta\bm{s}$ and residual hidden feature $\Delta\bm{h}$. The final outputs are the summation of these two branches
@@ 3.3.2 Uncertainty Estimation Module
where $\mathcal{R}$ is the set of rays in each batch. $\bm{n}(\bm{r})$ and $\hat{\bm{n}}(\bm{r})$ are the rendered normal and the pseudo ground-truth normal. Specifically, for the uncertainty value corresponding to a ray $\bm{r}$, we normalize it to $[0,1]$, denoted as $\bm{u}(\bm{r})$. When $\mathcal{L}_{prior}$ supervises the implicit SDF field together with the typical color loss, our proposed uncertainty values can be viewed as a trade-off between the color information and normal information. For complicated regions with high uncertainty, the proposed loss reduces the weight of normal supervision, thereby strengthening the role of color supervision. By balancing the supervision weights, our model can resist being misled by unreliable normal priors and effectively utilize more reliable normal priors.
To predict the pixel-wise uncertainty of estimated normal maps, we design a novel module shown in Figure 4. This module adopts a U-Net [46] architecture with RGB images and the estimated normal maps as input. To facilitate the uncertainty prediction, we also input high-level DINO [20] features, since the normal uncertainty is usually structure-relevant. We choose the visual foundation model DINO rather than CLIP [47], as DINO tends to capture pixel-level structure-aware image features during its self-supervised learning process, while CLIP features tend to be global and linguistic-related. Furthermore, a previous work [48] also demonstrates the effectiveness of DINO features in facilitating downstream visual understanding tasks, outperforming other foundation models (e.g., CLIP [47] and DeiT [49]). We extract features from the last attention layer of DINO, and then apply PCA technique to retain the most essential information. We train this uncertainty estimation module on a subset (disjoint from the test set) of the ScanNet [27]. Concretely, we compute the difference between the estimated normals and normals rendered from real 3D meshes to obtain the ground-truth uncertainty maps, and use them to supervise our proposed estimation module.
@@ 3.4 Loss Functions
The overall loss is the weighted sum of normal prior loss $\mathcal{L}_{prior}$, Eikonal loss $\mathcal{L}_{eik}$, and RGB color loss $\mathcal{L}_{rgb}$:
@@ 4.3.1 Analysis of Geometry Representation
Figure 13: Results of the novel view synthesis task. The first row is the ground-truth images, the second and third rows are novel views generated by volume rendering and rasterized by our reconstructed colored mesh, respectively.
@@ 4.3.2 Analysis of Appearance Embedding
Since we only supervise our model on 2D image observations, the inconsistency caused by photometric and environmental variations can lead to noisy geometries. Our experiments illustrate that the appearance embeddings can alleviate this problem, as the embeddings can learn to compensate for variations and guarantee a consistent reconstruction. We render the reconstructed normal maps and depth maps in Figure 10. The depth maps and normal maps are sharper and cleaner by adding appearance embeddings.
@@ 4.3.3 Analysis of Prior Enhancement
As illustrated in Figure 10, imperfections like motion blur during the video capture process can cause degraded predictions of normal priors, which hinder the subsequent reconstruction of fine and accurate geometries. After applying our proposed enhancement, the image becomes sharper, and the predicted normal maps get better. As a result, Table V shows that the accuracy and completeness of reconstructed meshes are improved as well.
@@ 4.3.4 Analysis of Uncertainty Estimation
The estimated uncertainty can serve as a measure of reliability and a balance between color supervision and normal supervision. In intricate areas, it guides our network to learn better geometry from rich color information rather than inaccurate normal information, thereby facilitating the correct reconstruction of exquisite structures. Ablation results show that the uncertainty estimation module helps to improve geometry quality. As fine-grained details take up only a small proportion of the scene and cannot be fully reflected by numerical metrics, we also visualize the reconstructed meshes and rendered normal maps in Figure 11. Adding uncertainty as guidance makes the reconstruction results more visually appealing and richer in high-frequency details. In addition, we find that supplementing the uncertainty estimation module with high-level DINO features can further facilitate the reconstruction of complex and delicate structures (shown in Figure 14).
@@ 4.5 Novel View Synthesis
The novel view is another form of understanding 3D scenes apart from meshes. Figure 13 shows the novel views generated by volume rendering, rasterized by our reconstructed colored mesh, and the corresponding ground-truth images. Note that while our primary goal is to reconstruct high-fidelity, detailed surfaces of an indoor scenario, we can also synthesize realistic images under novel views facilitated by the accurate reconstruction results.
@@ 4.6 Efficiency and Robustness
In Table VI, we compare the resource consumption of several baseline methods. ManhattanSDF [6] and MonoSDF [7] adopt the parameter compact MLP models but struggle to reconstruct fine geometries. The memory consumption and parameters of NeuralAngelo [26] are large because it utilizes 3D hash grids, which can be regarded as a special version of voxel grids. HelixSurf [35] represents the implicit SDF field with a relatively small MLP accompanied by a PatchMatch-based MVS module. However, the joint optimization of MVS module and neural radiance field in HelixSurf consumes a large memory overhead. In contrast, our method can produce the best quality meshes with comparable parameters and FLOPs, and less memory consumption.
@@ 5 Conclusion and Limitation
Acknowledgments
This work was supported by Natural Science Foundation of China (62332019, U2336214, 62202257), and The Talent Fund of Beijing Jiaotong University (2023XKRC045).